# Optimizing Probabilistic Conformal Prediction with Vectorized Non-Conformity Scores

### v1

@@ Abstract
Generative models have shown significant promise in critical domains such as medical diagnosis, autonomous driving, and climate science, where reliable decision-making hinges on accurate uncertainty quantification. While probabilistic conformal prediction (PCP) offers a powerful framework for this purpose, its coverage efficiency – the size of the uncertainty set – is limited when dealing with complex underlying distributions and a finite number of generated samples. In this paper, we propose a novel PCP framework that enhances efficiency by first vectorizing the non-conformity scores with ranked samples and then optimizing the shape of the prediction set by varying the quantiles for samples at the same rank. Our method delivers valid coverage while producing discontinuous and more efficient prediction sets, making it particularly suited for high-stakes applications. We demonstrate the effectiveness of our approach through experiments on both synthetic and real-world datasets.
@@ 1 Introduction
Uncertainty quantification is a critical component in predictive modeling and decision-making processes across a wide range of fields, including finance (Fujimoto et al.,, 2022; Blasco et al.,, 2024), healthcare (Seoni et al.,, 2023), and engineering (Kumar et al.,, 2023). In real-world applications, models often encounter inherent uncertainties due to data noise, model limitations, or unforeseen events. Accurately quantifying these uncertainties enables practitioners to assess risks and make informed decisions. For instance, in medical diagnostics, providing uncertainty estimates alongside disease predictions can guide clinicians in determining the necessity of further tests or interventions (Zou et al.,, 2023; Lambert et al.,, 2024). Similarly, in autonomous driving, understanding the uncertainty in object detection can enhance safety by prompting caution in ambiguous situations Michelmore et al., (2020); Su et al., (2023); Grewal et al., (2024).
Conformal prediction (Vovk et al.,, 2005) has emerged as a powerful framework for uncertainty quantification, offering statistically valid prediction sets without making distributional assumptions. Its popularity stems from its ability to provide finite-sample guarantees under the assumption of data exchangeability, making it widely applicable in various machine learning tasks. For example, in regression analysis, conformal prediction can generate prediction intervals that adapt to the local variability of the data, thereby offering more reliable uncertainty estimates compared to traditional methods.
Recent advancements, such as probabilistic conformal prediction (PCP) using conditional random samples (Wang et al.,, 2023), have paved the way for more effective uncertainty quantification in probabilistic and generative models. This approach leverages conditional sampling to construct the prediction sets in the form of a union of uncertainty balls around each generated sample, thus providing a discontinuous uncertainty set. This is a fundamental step towards better coverage efficiency (small set size), as ideally, we want the prediction set to avoid uncertain low-density regions.
However, the PCP method faces two major limitations: ($i$) When the number of generated samples is small, PCP may produce overly conservative prediction sets, as it relies on the minimal distance between the true response and generated samples. ($ii$) PCP treats all generated samples equally by using equal-sized balls, which lacks the flexibility to capture complex distributional dynamics. This uniform treatment can result in inefficiencies – producing large uncertainty sets – by not fully exploiting regions where the model is more confident.
To tackle these challenges, we propose a novel Probabilistic Conformal Prediction framework by Vectorizing the non-Conformity score with Ranked samples (PCP-VCR). Specifically, our approach introduces a non-conformity score vector that ranks generated samples based on their empirical density, capturing detailed information about the error distribution across samples. By individually optimizing the quantiles for samples at the same rank, we construct uncertainty sets that are more flexible and informative. The key idea is that by leveraging the ranked structure of the non-conformity scores, our method can allocate uncertainty in a more nuanced manner. High-ranking samples from dense regions, where the model is more confident, contribute to larger prediction sets, reflecting lower uncertainty. Conversely, lower-ranking samples from sparse regions may have smaller prediction sets or be excluded altogether due to higher uncertainty. This prioritization results in tighter prediction sets that are both flexible and efficient, providing practitioners with more informative uncertainty quantification. In our numerical study, we demonstrate our method shows promising empirical performance on both synthetic and real-world datasets, compared with state-of-the-art methods. In summary, the main contributions of this paper are:
@@
  • PCP-VCR: We present a novel method that utilizes vectorized non-conformity scores with ranked samples to enhance the efficiency of probabilistic conformal prediction.
@@
  • Theoretical and Empirical Validation: We demonstrate through theoretical analysis and empirical evaluations that PCP-VCR maintains valid coverage while significantly improving prediction set efficiency compared to baseline methods.
@@
In terms of conformal prediction, extended works have been done focusing on CP for deterministic model. Different summary statistics of the conditional distribution have been studied, for example mean (Shafer and Vovk,, 2008; Lei et al.,, 2018) and quantiles (Romano et al.,, 2019; Kivaranovic et al.,, 2020; Alaa et al.,, 2023). In Sesia and Romano, (2021), the authors use histograms to estimate the conditional distribution and provide efficient prediction interval with approximate conditional coverage. Though the above methods all provide continuous prediction intervals, they cannot adapt to multimodality well. In Lei et al., (2013) and Lei and Wasserman, (2014), the authors first consider the aspect of density level sets and density-based clustering for providing efficient prediction set. The authors use kernel density estimator to construct the non-conformity scores and with a proper choice of bandwidth, the prediction set could be discontinuous. It is also proved that the method is asymptotically near-optimal when the number of calibration data goes to infinity. There are also some other works (Izbicki et al.,, 2019; Chernozhukov et al.,, 2021; Izbicki et al.,, 2022) that estimate the density or cumulative distribution function to construct prediction set. As a reminder, in our setting, we only need access to the random samples generated from the fitted model instead of explicit density estimation. So our setting is more generic and applicable to real applications. In Tumu et al., (2024), the authors fit shape functions (convex hull, hyperrectangle, and ellipsoid) on the calibration data and minimize the total volume of shapes to reduce the inefficiency. In Guha et al., (2024), the authors convert regression to a classification problem and can provide discontinuous prediction set with interpolation of classification scores.
@@
Efficiency of the CP prediction set is a fundamental assessment of the quality of CP method. Some works explore the effect of different non-conformity score measures such as by scaling and reweighting the non-conformity score of each calibration data (Papadopoulos et al.,, 2002, 2011; Kivaranovic et al.,, 2020; Bellotti,, 2020; Amoukou and Brunel,, 2023). Specifically, in H. Zargarbashi et al., (2023), the authors utilize the network structure to construct a non-conformity score that incorporates neighborhood information. On the other hand, some works (Stutz et al.,, 2021; Einbinder et al.,, 2022) directly reduce inefficiency by incorporating a differentiable inefficiency measurement into the training loss. Bai et al., (2022) reformulate the conformal prediction as a constrained optimization problem. By introducing an extra parameter and enriching the search space of the optimization problem, it demonstrates better performance in efficiency. Our main idea is partially motivated by this optimization reformulation. We also gain insights from the recent development of CP for multi-target regression tasks and the definition of multivariate quantile function as discussed in (Messoudi et al.,, 2021; Sun and Yu,, 2022).
@@ 2 Preliminaries
In this section, we provide an overview of Conformal Prediction and Probabilistic Conformal Prediction (Wang et al.,, 2023), respectively, along with the key notations that will be used throughout the paper.
2.1 Conformal Prediction
@@
This ensures that the constructed prediction set $\hat{\mathcal{C}}(\cdot)$ covers the true label with a probability of at least $1-\alpha$, where the probability is marginal over the calibration data and the test point.
To achieve this, CP relies on computing the non-conformity score, which measures the extent to which a new data point conforms to the behavior of the training data under a predictive model. For regression tasks, a typical choice of non-conformity score is the absolute residuals from a predictive model $\hat{f}$, i.e., $E_{i}=\|Y_{i}-\hat{f}(X_{i})\|$. Based on this score, the empirical quantile of the non-conformity scores is calculated, and the prediction set $\hat{\mathcal{C}}(X_{n+1})$ is constructed such that the miscoverage rate condition is satisfied.
@@
2.2 Probabilistic Conformal Prediction
Probabilistic Conformal Prediction (PCP) (Wang et al.,, 2023) extends the CP framework to probabilistic (or generative) models. Given a fitted conditional probabilistic model $\hat{p}(Y|X)$, the approach involves sampling from the model rather than making a point prediction. Specifically, for each calibration point $(X_{i},Y_{i})$, we generate $K$ random samples
${\hat{Y}_{i,1},\dots,\hat{Y}_{i,K}}\sim\hat{p}(Y|X_{i}).$
The non-conformity score is then defined as the minimum distance between the true response $Y_{i}$ and the generated samples:
$E_{i}=\min_{k=1,\dots,K}\|Y_{i}-\hat{Y}_{i,k}\|.$ (2)
For a new test point $(X_{n+1},Y_{n+1})$, the generative model produces $K$ random samples ${\hat{Y}_{1},\dots,\hat{Y}_{k}}$. The prediction set is constructed as the union of $K$ regions:
$\hat{\mathcal{C}}(X_{n+1})=\cup_{k=1}^{K}\hat{\mathcal{C}}_{k},$ (3)
where each region $\hat{\mathcal{C}}_{k}$ is defined by
Here $Q(\cdot;\alpha)$ is the $1-\alpha$ quantile of the empirical distribution of the non-conformity scores $E_{1},\dots,E_{n}$. If the error measure used is the $\ell_{2}$ norm, the resulting prediction set is a union of balls centered at each generated sample $\hat{Y}_{k}$, with a constant radius determined by the quantile $Q(\alpha)$.
A significant limitation of the PCP is its inefficiency, especially when $K$ is limited. The non-conformity score in (2) relies on the minimal distance between the true response and all generated samples, resulting in prediction sets that are unions of equal-sized balls centered at the generated samples. If $K$ is small, there’s a high chance that none of the generated samples $\{\hat{Y}_{i,k}\}_{k=1}^{K}$ are close to $Y_{i}$, leading to large nonconformity scores $E_{i}$. As a result, the quantile $Q(\alpha)$ becomes large to maintain the desired coverage level, causing the prediction sets $\hat{\mathcal{C}}(X_{n+1})$ to expand significantly and reducing their efficiency.
Figure 1: An illustrative comparison betwen PCP and proposed PCP-VCR, where PCP-VCR provides a more flexible and efficient prediction set. The non-conformity score vector $(E_{i,1},\dots,E_{i,K})^{\top}$ provides more detailed information on the underlying distribution represented by the blue area. Samples in denser regions (darker blue) are ranked higher (small $r$), indicating greater model confidence and larger $\mathcal{C}_{r}$ sizes, while low-rank scores (large $r$) capture uncertainty in sparser regions, where $\mathcal{C}_{r}$ sizes are smaller.
@@ 3 Proposed Method
In this section, we introduce a novel method for probabilistic conformal prediction, which vectorizes the non-conformity score with ranked samples, referred to as PCP-VCR.
The key idea of our method is to introduce a vectorized non-conformity score, which ranks the generated samples based on their empirical density. Unlike traditional non-conformity measures, such as the $\ell_{2}$ norm, which yield a single scalar value, our vectorized approach captures more detailed information about the distribution of errors across samples. We then jointly optimize the quantile for each ranked sample, maximizing the efficiency of the uncertainty set while preserving the validity guarantee. The uncertainty set is constructed by placing an uncertainty ball around each sample, with the radius of each ball determined by its respective optimized quantile, leading to more flexible and informative uncertainty quantification. Figure 1 illustrates a comparison between our method and the traditional PCP approach. The detailed steps of our proposed method are outlined in Algorithm 1.
Algorithm 1 PCP-VCR
@@ 3.1 Vectorized Non-Conformity Score
The rationale of constructing the non-conformity score vector with ranked samples are two-fold: ($i$) By forming a vector of non-conformity scores rather than a single scalar value, we capture more detailed information about the relationship between the true label $Y_{i}$ and the generated samples. Each element $E_{i,r}$ in the vector corresponds to a specific rank $r$, reflecting different aspects of the sample distribution. This enriched information enables us to construct more flexible and informative prediction sets compared to traditional CP/PCP methods that rely on a scalar score. ($ii$) Ranking the samples based on their average $m$-nearest neighbor distances allows us to identify and prioritize samples located in denser regions of the estimated model $\hat{p}$. Samples with smaller $\bar{D}_{i,j(1)}$ are presumed to be in high-density areas and are thus ranked higher. Consequently, the non-conformity scores $E_{i,r}$ for different ranks $r$ carry varying significance: High-rank scores ($r$ is small) correspond to samples in denser regions, where the model is more confident, whereas low-rank scores ($r$ is large) represent samples in sparser regions, capturing the model’s uncertainty. By comparing non-conformity scores across different calibration data points at the same rank, we can assess discrepancies at multiple levels of the distribution. This stratified approach provides a more nuanced understanding of the model’s performance across the entire output space, leading to more accurate and adaptive uncertainty sets.
@@ 3.2 Efficiency-Enhanced Prediction Set
The prediction set can be constructed as follows: First, we compute the non-conformity score vectors $\{E_{i}\coloneqq(E_{i,1},\dots,E_{i,K})^{\top}\}_{i=1}^{n}$ for the calibration data $\mathcal{D}$. Then, we define the prediction set as follows:
@@
and $Q_{r}(\cdot;\beta_{r})$ is the $1-\beta_{r}$ quantile of the empirical distribution of the $r$-th ranked non-conformity scores across $n$ samples, $E_{1,r},\dots,E_{n,r}$. For notational simplicity, we use $Q_{r}(\beta_{r})$ throughout the paper.
Unlike PCP, which constructs its prediction set using equal-sized balls as defined in (3), the radius of each ball $\hat{\mathcal{C}}_{r}(\beta_{r})$ in (6) is rank-dependent, determined by $Q_{r}(\beta_{r})$. By introducing the quantile vector $(Q_{r}(\beta_{1}),\dots,Q_{r}(\beta_{K}))^{\top}$, we allow each rank $r$ to have its own quantile level $\beta_{r}$, providing greater flexibility in constructing the prediction set. This approach enables us to optimize the quantiles individually for each rank, leading to more efficient and tighter uncertainty sets while still ensuring valid coverage – even though the $\beta_{r}$ values are not necessarily equal to the user-defined miscoverage rate $\alpha$.
@@
where $d$ is the dimensionality of the response variable $Y$. The coverage rate can be estimated by
$\frac{1}{n+1}\sum_{i=1}^{n}\hat{S}_{i}(\beta),$ (7)
where $\hat{S}_{i}(\beta)$ equals $1$ if there is at least one ball in the prediction set covers the calibration data $Y_{i}$, i.e.,
@@
Formally, given the non-conformity score vectors, we formulate the following optimization problem:
$\min_{\beta}\left\{\sum_{r=1}^{K}(Q_{r}(\beta_{r}))^{d}\right\},~{}\text{s.t.}~{}\frac{1}{n+1}\sum_{i=1}^{n}\hat{S}_{i}(\beta)\geq 1-\alpha.$ (8)
@@
We note that the proof of Theorem 3.1 is non-trivial. Compared to other conformal prediction validity theorems, which typically deal with scalar non-conformity scores, the novelty here is that it extends the coverage validity to a more complex setting of vectorized and ranked non-conformity scores. For scalar non-conformity scores, the quantile is unique and can be readily determined by sorting the scalar values. However, with multivariate vectors, two primary challenges arise in computing the quantile: ($i$) The quantile is represented by a vector, in which its value is not unique ($ii$) There is no (total) ordering between vectors, meaning that vectors are not fully comparable and thus cannot be sorted as scalars are. The theorem addresses these challenges of computing valid quantiles for vectors. By defining a partial ordering between vectors and optimizing across this structure, the theorem ensures that even with this increased complexity, the prediction set maintains the desired level of validity while offering more efficient and adaptive uncertainty quantification.

### v2

# Generative Conformal Prediction with Vectorized Non-Conformity Scores

@@ Abstract
Conformal prediction (CP) provides model-agnostic uncertainty quantification with guaranteed coverage, but conventional methods often produce overly conservative uncertainty sets, especially in multi-dimensional settings. This limitation arises from simplistic non-conformity scores that rely solely on prediction error, failing to capture the prediction error distribution’s complexity. To address this, we propose a generative conformal prediction framework with vectorized non-conformity scores, leveraging a generative model to sample multiple predictions from the fitted data distribution. By computing non-conformity scores across these samples and estimating empirical quantiles at different density levels, we construct adaptive uncertainty sets using density-ranked uncertainty balls. This approach enables more precise uncertainty allocation – yielding larger prediction sets in high-confidence regions and smaller or excluded sets in low-confidence regions – enhancing both flexibility and efficiency. We establish theoretical guarantees for statistical validity and demonstrate through extensive numerical experiments that our method outperforms state-of-the-art techniques on synthetic and real-world datasets.
@@ 1 Introduction
Conformal prediction (CP) (Vovk et al.,, 2005) has gained prominence as a robust framework for uncertainty quantification (UQ), offering prediction sets with guaranteed coverage under minimal distributional assumptions. As a model-agnostic approach, CP can enhance any trained machine learning models to output not just point estimates but also provide uncertainty sets that contain the unobserved ground truth with user-specified high probability. As a result, CP has found broad applications across fields such as healthcare (Seoni et al.,, 2023), finance (Fujimoto et al.,, 2022; Blasco et al.,, 2024), and autonomous systems (Michelmore et al.,, 2020; Su et al.,, 2023; Grewal et al.,, 2024), where reliable decision-making under uncertainty is critical. By delivering user-specified confidence levels, CP has become an indispensable tool for practitioners striving to connect predictive modeling with actionable, real-world insights.
While CP has proven to be a reliable and versatile framework, its prediction sets can sometimes be overly conservative (Messoudi et al.,, 2021), limiting their practicality and interpretability in some real-world applications. This issue often stems from the simplifying assumptions used in defining the non-conformity scores, such as the common adoption of the $\ell_{2}$ distance. While effective in one-dimensional settings, this definition often fails to account for the complexity of multi-dimensional data distributions, especially when the data resides on a manifold (Kuleshov et al.,, 2018).
Several approaches have been proposed to address these limitations. For example, instead of computing the $\ell_{2}$ distance between vectors, Neeven and Smirnov, (2018) considered using elementwise distances and stacked multiple regression models to provide the uncertainty set, though it lacks theoretical validity and does not account for the dependency between variables. Other works (Messoudi et al.,, 2021; Sun and Yu,, 2024) integrate copula function with conformal prediction to incorporate the variable dependency and provide theoretical guarantees. Yet, these methods are specifically designed for multi-target regression tasks, limiting their applicability to broader predictive tasks. More recent efforts aim to refine non-conformity scores to better capture prediction errors and adapt the prediction set to each variable. For instance, covariance-weighted non-conformity scores (Johnstone and Cox,, 2021; Messoudi et al.,, 2022; Xu et al.,, 2024) provide ellipsoidal uncertainty regions that can be tailored to each data dimension. However, a fundamental limitation persists: these methods rely solely on model prediction errors and fail to leverage additional information about the model’s predictive distribution, which could further enhance uncertainty estimation.
To address the limitations outlined above, in this paper, we introduce a novel generative conformal prediction framework with vectorized non-conformity scores to improve prediction efficiency by capturing richer distributional information. Instead of relying on a predictive model, we train a generative model to sample multiple predictions from the target distribution. Non-conformity scores are computed across these samples, and empirical quantiles are estimated at different density levels, respectively. Uncertainty sets are then constructed using adaptive-radius balls centered at the generated samples, with radii determined by the estimated quantiles. By optimally selecting quantiles across different density levels, our method produces more flexible and efficient uncertainty sets. Such density-ranked approach enables a more nuanced allocation of uncertainty: high-ranking samples from dense regions – where the model exhibits greater confidence – contribute to larger prediction sets, reflecting lower uncertainty, while lower-ranking samples from sparse regions are associated with smaller or excluded prediction sets, indicating higher uncertainty. We also establish theoretical guarantees on the statistical validity of the prediction sets. Extensive numerical studies demonstrate that our approach outperforms state-of-the-art methods on both synthetic and real-world datasets. In summary, the main contributions of this paper include:
@@
  • GCP-VCR: We present a novel method utilizing vectorized non-conformity scores with ranked samples to enhance the efficiency of generative CP.
@@
  • Theoretical and Empirical Validation: We demonstrate through theoretical analysis and empirical evaluations that GCP-VCR maintains valid coverage while significantly improving prediction set efficiency compared to baseline methods.
@@
In terms of conformal prediction, extended works have been done focusing on CP for deterministic models. Different summary statistics of the conditional distribution have been studied, for example, mean (Shafer and Vovk,, 2008; Lei et al.,, 2018) and quantiles (Romano et al.,, 2019; Kivaranovic et al.,, 2020; Alaa et al.,, 2023). In Sesia and Romano, (2021), the authors use histograms to estimate the conditional distribution and provide efficient prediction intervals with approximate conditional coverage. Though the above methods all provide continuous prediction intervals, they cannot adapt to multimodality well. In Lei et al., (2013) and Lei and Wasserman, (2014), the authors first consider the aspect of density level sets and density-based clustering for providing efficient prediction set. The authors use kernel density estimator to construct the non-conformity scores and with a proper choice of bandwidth, the prediction set could be discontinuous. There are also following works (Izbicki et al.,, 2019; Chernozhukov et al.,, 2021; Izbicki et al.,, 2022) that estimate the density or cumulative distribution function to construct prediction sets. As a reminder, in our setting, we only need access to the random samples generated from the fitted model instead of explicit density estimation. So our setting is more generic and applicable to real applications. Other than density estimation, (Tumu et al.,, 2024) fits shape functions (convex hull, hyperrectangle, and ellipsoid) on the calibration data to provide a discontinuous prediction set as a union of fitted shapes. In Guha et al., (2024), the authors convert regression to a classification problem and can provide discontinuous prediction sets with interpolation of classification scores.
@@
Efficiency, measured by the size of the prediction set, is a fundamental assessment of the quality of a CP method. Some works explore the effect of different non-conformity score measures such as by scaling and reweighting the non-conformity score of each calibration data (Papadopoulos et al.,, 2002, 2011; Kivaranovic et al.,, 2020; Bellotti,, 2020; Amoukou and Brunel,, 2023). On the other hand, some works (Stutz et al.,, 2021; Einbinder et al.,, 2022) directly reduce inefficiency by incorporating a differentiable inefficiency measurement into the training loss. Bai et al., (2022) reformulate CP as a constrained optimization problem. By introducing an extra parameter and enriching the search space of the optimization problem, it demonstrates better performance in efficiency. Our main idea is partially motivated by this reformulation. We also gain insights from the recent development of CP for multi-target regression tasks and the definition of multivariate quantile function as discussed in (Messoudi et al.,, 2021; Sun and Yu,, 2024).
Recent advancements, such as probabilistic conformal prediction (PCP) using conditional random samples (Wang et al.,, 2023), have paved the way for CP in probabilistic and generative models. PCP leverages generated random samples to construct the prediction sets in the form of a union of balls around each generated sample, offering a discontinuous uncertainty set. This is a fundamental step towards better coverage efficiency, as ideally, we want the prediction set to avoid low-density regions. However, PCP faces two major limitations: i) PCP does not distinguish different generated samples and assigns the same radius to each ball, ii) when the number of generated samples is relatively small, PCP can yield overly conservative prediction sets. We will discuss the comparison of our proposed method and PCP in details in section 3.3.
@@ 2 Preliminaries
This ensures that the prediction set $\hat{\mathcal{C}}(\cdot)$ covers the true label with a probability of at least $1-\alpha$, where the probability is marginal over the calibration data and the test point.
To achieve this, CP computes a non-conformity score, which measures the extent to which a new data point conforms to the behavior of the training data under a predictive model. For regression tasks, a typical choice of non-conformity score is the absolute residuals from a predictive model $\hat{f}$, i.e., $E_{i}=|Y_{i}-\hat{f}(X_{i})|$. Based on this score, the empirical quantile of the non-conformity scores of calibration data is calculated to construct prediction set $\hat{\mathcal{C}}(X_{n+1})$ such that the miscoverage rate condition (1) is satisfied.
@@
Figure 1: An illustrative comparison betwen the proposed GCP-VCR and PCP, where GCP-VCR provides a more flexible and efficient prediction set. The non-conformity score vector $(E_{i,1},\dots,E_{i,K})^{\top}$ provides more detailed information on the underlying distribution represented by the blue area. Samples in denser regions (darker blue) are ranked higher (small $r$), indicating greater model confidence and larger $\mathcal{C}_{r}$ sizes, while low-rank scores (large $r$) capture uncertainty in sparser regions, where $\mathcal{C}_{r}$ sizes are smaller.
@@ 3 Proposed Method
In this section, we introduce a novel method, referred to as Generative Conformal Prediction with Vectorized Non-Conformity Scores (GCP-VCR), which utilizes a vectorized non-conformity score with ranked samples.
Generative models offer a versatile framework for capturing complex data distributions, advancing beyond deterministic models that output limited summary statistics (e.g., mean or quantiles). Let $\hat{p}(Y|X)$ denote a fitted generative model from which conditional random samples can be drawn, e.g., $\hat{Y}_{1},\dots,\hat{Y}_{K}\sim\hat{p}(Y|X)$. These random samples encode nuanced information of the prediction error distribution—such as multimodal patterns and tail behaviors—that traditional scalar non-conformity scores fail to capture. Leveraging this information enables more robust and efficient uncertainty quantification. To fully exploit this information, we propose a vectorized non-conformity score that aggregates prediction errors across all generated samples. We further rank these samples by their empirical density, prioritizing high-density regions where the model is confident and assigning adaptive quantile levels to samples at different ranks. This ranking mechanism enables localized uncertainty allocation by distinguishing high-confidence and low-confidence samples. We formalize this process as a constrained optimization problem to derive the optimal quantile vector, ensuring both statistical validity and uncertainty set efficiency. Figure 1 highlights our advancements compared to PCP, while Algorithm 1 outlines the complete workflow. The following sections detail our method step by step rigorously.
Algorithm 1 GCP-VCR
@@ 3.1 Vectorized Non-Conformity Score
The rationale of constructing the non-conformity score vector with ranked samples is two-fold: ($i$) By forming a vector of non-conformity scores rather than a single scalar value, we capture more detailed information about the relationship between the true label $Y_{i}$ and the generated samples. Each element $E_{i,r}$ in the vector corresponds to a specific rank $r$, reflecting different aspects of the sample distribution. This enriched information enables us to construct more flexible and informative prediction sets compared to traditional CP methods that rely on a scalar score. ($ii$) Ranking the samples based on their average $m$-nearest neighbor distances allows us to identify and prioritize samples located in denser regions of the estimated model $\hat{p}$. Samples with smaller $\bar{D}_{i,j(1)}$ are presumed to be in high-density areas and are thus ranked higher. Consequently, the non-conformity scores $E_{i,r}$ for different ranks $r$ carry varying significance: High-rank scores ($r$ is small) correspond to samples in denser regions, where the model is more confident, whereas low-rank scores ($r$ is large) represent samples in sparser regions, capturing the model’s uncertainty. By comparing non-conformity scores across different calibration data points at the same rank, we can assess discrepancies at multiple levels of the distribution. This stratified approach provides a more nuanced understanding of the model’s performance across the entire output space, leading to more accurate and adaptive uncertainty sets.
We note that the choice of distance measure in (3) can be adapted based on the nature of $\mathcal{Y}$. For instance, if $\mathcal{Y}$ resides in Euclidean space, the $\ell_{p}$ norm may be adopted. In the context of geographical data, the distance measure could be the geodesic distance or even a general kernel function, depending on the application.
@@ 3.2 Efficiency-Enhanced Prediction Set
The prediction set can be constructed as follows: First, we compute the non-conformity score vectors $\{E_{i}\coloneqq(E_{i,1},\dots,E_{i,K})^{\top}\}_{i=1}^{n}$ for the calibration data $\mathcal{D}$. Then, for a new test pair $(X_{n+1},Y_{n+1})$, we define the prediction set as follows:
@@
and $Q_{r}(\cdot;\beta_{r})$ is the $(1-\beta_{r})$-th quantile function. For notational simplicity, we use $Q_{r}(\beta_{r})=Q_{r}(\{E_{1,r},\dots,E_{n,r}\}\cup\{\infty\};\beta_{r})$ throughout the paper.
If the distance measure in (3) is the $\ell_{2}$-norm, the resulting prediction set becomes a union of balls centered at each generated sample $\hat{Y}_{n+1,j(r)}$, with rank-dependent radius determined by $Q_{r}(\beta_{r})$. Introducing the quantile vector $(Q_{r}(\beta_{1}),\dots,Q_{r}(\beta_{K}))^{\top}$ allows each rank $r$ to have its own quantile level $\beta_{r}$, offering greater flexibility in constructing prediction sets. This approach enables us to optimize the quantiles individually for each rank, leading to more efficient and tighter uncertainty sets while still ensuring valid coverage – even though the $\beta_{r}$ values are not necessarily equal to the user-defined miscoverage rate $\alpha$. As a comparison, PCP provides the uncertainty set as a union of equal-sized balls; a detailed comparison with PCP will be presented in Section 3.3.
@@
where $d$ is the dimensionality of the response variable $Y$. The calibration coverage rate is defined as
$\frac{1}{n+1}\sum\limits_{i=\{1,\dots,n,\infty\}}\hat{S}_{i}(\beta),$ (5)
where we let $E_{\infty,r}=\infty$ for any $r$, and define
@@
Intuitively, each $\hat{S}_{i}(\beta)$ equals $1$ if there is at least one ball in the prediction set covers the calibration data $Y_{i}$. Here, the denominator $n+1$ and use of $\hat{S}_{{\infty}}(\beta)$ is for theoretical proof of coverage validity, and can be understood as a finite-sample correction. Formally, given the non-conformity score vectors, we formulate the following optimization problem:
$\begin{gathered}\min_{\beta}\left\{\sum_{r=1}^{K}(Q_{r}(\beta_{r}))^{d}\right\},\\ ~{}\text{s.t.}~{}\frac{1}{n+1}\sum\limits_{i=\{1,\dots,n,\infty\}}\hat{S}_{i}(\beta)\geq 1-\alpha.\end{gathered}$ (6)
@@
We note that the proof of Theorem 3.1 is non-trivial, as it involves handling non-conformity score vectors. For scalar non-conformity scores, the quantile is unique and can be readily determined by sorting the scalar values. However, with multivariate vectors, two primary challenges arise in computing the quantile: ($i$) The quantile is represented by a vector, in which its value is not unique ($ii$) There is no (total) ordering between vectors, meaning that vectors are not fully comparable and thus cannot be sorted as scalars are. To overcome these challenges, we define a partial ordering between vectors and select the vector that maximizes efficiency while ensuring validity.